a block measures mass 50 cm by 20 cm by 5 cm and has a mass of 200 g , calculate the maximum and minimum pressure
1
Expert's answer
2020-06-10T18:25:03-0400
"P_{max}=\frac{mg}{bc}=\frac{(0.2)(9.8)}{(0.2)(0.05)}=200\ Pa"

"P_{min}=\frac{mg}{ba}=\frac{(0.2)(9.8)}{(0.2)(0.5)}=20\ Pa"